# Differentiable Physics-based System Identification for Robotic Manipulation of Elastoplastic Materials

### v1

@@ 3.4 Real-to-sim object reconstruction
Real-world platform: To collect real-world data, we set up a deformable object manipulation system where a Kuka IIWA LBR 14 industrial arm Kuka (2024) is equipped with a Zivid One+ medium camera Zivid (2024) for perceiving the object, and one of the three end-effectors, namely a rectangular cuboid, a cylinder roller and a bullet-shaped object for collecting interaction experiences with different contact geometries, as shown at the bottom-left of Figure 1. We use plasticine (non-hardening modelling clay) as the main material for our experiments.
@@
Reconstruction pipeline: To simulate the object as a set of particles, we design a pipeline to reconstruct the object particle system $\mathcal{X}^{rec}$ from the given fusion point cloud, as shown by the orange boxes in Figure 2. Firstly, we down-sample the processed fusion point cloud and create a watertight mesh using the ball-pivoting (BP) algorithm Bernardini et al. (1999). The artificial bottom surface points prevent the BP algorithm from creating a non-flat bottom mesh surface. To sample particles, we create a set of voxels that fill in the reconstructed mesh surface uniformly given a resolution ($1080\times 1080$). Lastly, we fill particles into the voxels with a given density ($4\times 10^{6}$ particles$/m^{3}$). We assume that the particles are uniformly distributed within the mesh.
@@
Table 1: Waypoint designs and statistics of the interaction motions. For collecting optimisation data, there are two motions for each of the two levels of contact complexity. For out-of-distribution validation, one motion per end-effector is designed. Directions of the waypoints are relative to the robot base frame, as shown in Figure 1. Each waypoint of the real trajectories takes uneven time interval, while each waypoint in simulation takes exactly $0.01$ second.
@@ 3.5 Real-to-sim trajectory reconstruction
All motions start from a configuration where the end-effector tip is positioned at the top centre of the object. ROS Quigley et al. (2009) and the MoveIt!-based Görner et al. (2019) OMPL planner Sucan et al. (2012) are used to plan real-world motion trajectories $\tau^{real}$. For calculating each motion plan, we pass a series of waypoints to the MoveIt! planner by discretising each segment of the motion with an interval of $0.002$ m or $5^{\circ}$.
A challenging phenomenon that occurs during our data collection process is that the object tends to stick to the end-effectors after contact is made. If we allow this to happen and assume that the object always drops down eventually, the optimisation process will be extremely difficult due to a large uncertain dynamics process of the object dropping from the air. For example, the object may bounce out of the workspace or even off the table. This further exacerbates the difficulty of optimising with only the start and end observations of the object. However, it is beyond the capability of current perception hardware and thus a future research direction. To simplify the problem, we take a simple workaround by covering the end-effector with a thin layer of flour before each manipulation motion is executed. This greatly prevents the sticking phenomenon from happening. Note that the end-effector friction coefficient we are optimising for is then the one covered by flour instead of the original value.
Time duration constrained reconstruction: To simulate end-effector motions, the real-world trajectory generated by MoveIt! is inconvenient as it has uneven time differences between consecutive waypoints and the simulation can only handle a constant step size $dt$. Thus, for a motion segment between a pair of waypoints, we reconstruct its simulation counterpart to have constant velocity by dividing the travelled distance by the real time duration which is provided by the motion planner, and then we discretise the segment with a constant $dt$. In this study, we set $dt=0.01$ second for better simulation stability. A larger $dt$ will result in a too-high compounding error during simulation stepping, while a too small $dt$ will demand too much computation. The statistics of the real and reconstructed motions are summarised in Table 1. It can be seen that the validation motions are much longer than the optimisation ones. The MoveIt! trajectories and the reconstructed ones are saved as .npy files and will be open to the public upon acceptance.
@@ 3.6 Loss functions
Finally, although the CD and EMD losses are very common in calculating point set distance, they are not intuitive to visualise and understand. Also, as revealed by the experiment results shown in Figures 4 and  5, the CD and EMD losses focus on different spatial aspects of two point sets, which make them biased for result analysis. To make results easier to analyse and compare in an unbiased way, we further calculate the heightmaps of the fusion point cloud and the simulated particles, denoted as $\mathcal{I}_{T}^{real}$ and $\mathcal{I}_{T}^{sim}$. Specifically, we discretise a $0.11\times 0.11\text{m}^{2}$ square area centred at the centre of the object into a $32\times 32$ grid. For each cell of the grid, we find the point or particle in the x-y range with the highest $z$ value and assign that value as the height of that cell. To compare the distance between two heightmaps, we simply take their summed pixel difference as follows.
@@ 4.1 Experiment design
For each datapoint, the real-world object is initially roughly shaped into a convex shape. We then acquire its point cloud by fusing the captured multi-view data. Next, we calculate the coordinates of the top centre of the object, where the end-effector tip will be moved to, and the motion will always be executed from there. We perform the same point cloud capturing and fusion process after a motion is executed. Each datapoint takes about $2$ to $3$ minutes to collect, with most of the time spent on capturing point clouds.
Optimisation & validation: At each contact level, for each pair of loss function and dataset, gradient descent is carried out with the Adam algorithm Kingma and Ba (2015) with three random seeds for $100$ iterations (gradient updates). For each datapoint, the simulation loads the initial particle configuration, simulates the motion and produces the resultant particle state, which is used to compute losses and gradients. Each optimisation iteration goes over all datapoints within a dataset and takes the average gradients to update the parameters. In-distribution validation is done after every gradient update with the validation dataset, simulating all datapoints and calculating the losses.
Table 2 displays the step sizes and the value ranges of the parameters of interest for gradient updates. To determine the appropriate step sizes, we randomly sample $20$ sets of parameter values and compute the statistical means and standard deviations of the gradients for each parameter based on the 12-mix dataset. We then randomly initialise the parameters, run the Adam optimiser for $100$ gradient updates with uniformly randomly generated gradients using the statistical means and deviations with different step sizes, and select the step size that allows the randomised parameters to converge in about $70$ gradient updates. The value ranges of the parameters are selected according to either the definitions of the physics models Jones (2009); Stomakhin et al. (2013) or values reported by various research on manual parameter identification experiments with similar real-world materials, such as soil and clay StructX (2014-2024); Waheed and Asmael (2023). However, a too-high value of Young’s modulus $E$ was found to cause numerical instability in the simulation with the particles exploding away. Thus, an empirical upper bound for $E$ was determined by slowly increasing it until the simulation became unstable.
@@ 4.3 In-distribution performances
  • Does DPSI produce parameter values that are physically realistic and interpretable?
@@
However, the parameter distributions in Figure 5 clearly show that some parameters converge to much smaller and distinct value regions than those found in the level-1 experiments. More specifically, the values of Young’s modulus ($E$), Poisson’s ratio ($\nu$) and yield stress ($\boldsymbol{\sigma}_{y}$) become more converged, while they are more dispersed in the level-1 results. Interestingly, the material density ($\rho$) remains quite dispersed at both contact complexity levels. This is not surprising because, compared to the level-1 motions, the higher contact complexity induced by the shifting parts of the level-2 motions naturally poses a stricter selection range for the physics parameter values. Also, some parameter values are closer to empirical studies of soft/hard/saturated clay in the level-2 experiments. For example, the Young’s modulus values are closer to the reported range of $5000$ to $54000$ kPa, the material density seem to be closer to $1400$ kg/m3, and Poisson’s ratio are closer to the reported range of $0.4$ to $0.5$ Waheed and Asmael (2023); StructX (2014-2024). However, the reported yield stress values for clays ($210$ to $600$ kPa) Rehman et al. (2018) were found to make the material collapse in our simulator. A comparison of simulating the materials using the reported parameters in the literature and the ones found by DPSI is procided in the fifth subsections of the Results section.
@@
Table 3: The parameter values corresponding to the lowest validation heightmap distance in each optimisation case at both contact complexity levels. Bolded texts denote the three lowest losses at each level.
@@
Thirdly, the results show that it is possible to obtain comparable performances with just one datapoint at both contact complexity levels. This can be concluded by comparing the results horizontally: 1) most validation curves of the top five rows in Figs. 4 and 5 with 1-datapoint show highly similar tendencies with the 6-mix and 12-mix datasets, and 2) the last four rows in Figs. 4 and last six rows in Figure 5 also show that the parameters found with different datasets mostly converge to similar value regions. Moreover, Table 3 shows that the lowest validation heightmap distance at level-1 contact complexity and the third lowest at level-2 are achieved by optimising with the 1-rec. and 1-round datasets. These results are aligned with the analysis of the loss landscapes, where similar loss distributions are observed among different datasets.
Figure 6: Validation losses (top five rows) and parameter values (last four rows) over $100$ gradient updates at level-1 contact complexity from optimising the PRT EMD loss function with an extra set of $8$ random seeds. Each column presents the results of optimising with a different dataset. Each row shows the changes of an evaluation metric or a parameter, denoted on the left. In each figure, different colours indicate the results of a different random seed, as labelled by the legend on the top.
Lastly, we perform extra experiments of optimising the PCD EMD loss with the three 1-datapoint datasets to examine the random seed sensitivity of DPSI. The validation losses and parameters are shown in Figure 6. On the one hand, these runs all converge to stable (though different) loss and parameter values, still reveal the nagative relationship between the CD and EMD losses, and can achieve comparable performances with results from larger datasets. On the other hand, it can be seen that the main cause of converging to different solutions is the different initial parameter values, which were determined by the peusdo random process controlled by different random seed values. This is not suprising as gradient-descent algorithms with complex physics problems are known to be sensitive to the choice of the initial solution Hu et al. (2020); Antonova et al. (2023). Therefore, with a good initial guess of the solution, DPSI is likely to achieve desirable system identification results.
@@
Figure 7: Visualisations of two configurations of the simulated objects and their heightmaps after applying the second motions at contact level 1 (left) and 2 (right), simulated with the best set of physics parameters. Darker colours of the heightmaps indicate greater heights of the object. For each combination of the loss function and dataset at each level, three trajectories are simulated (3 effectors $\times$ 1 datapoints). The results for each trajectory are grouped with the ground truth placed on the left. In each group, a row shows the results corresponding to a loss function and a column results corresponding to a dataset.
@@
Qualitative results: We now examine the manipulation processes simulated using the best physics parameters corresponding to the lowest validation heightmap losses as shown in Table 3. Figure 7 shows the particles and heightmaps after the manipulation for two object configurations for each pair of contact complexity and end-effector.
@@
Secondly, by examining the details of the heightmaps more carefully, it shows that optimising the CD losses tends to produce heightmaps with darker colours (which represent greater heights of the object), while optimising the EMD losses tends to produce heightmaps where the objects look slightly bigger and wider. It indicates that optimising the CD losses tends to produce particles that match the height of the ground truth shapes, i.e., in the $z$ direction, while optimising the EMD losses tends to match the $x$ and $y$ directions. This is especially true at level-1 contact complexity. At level 2, the PRT CD loss sometimes (compare the second rows of the level-2 heightmaps with the rectangle and round end-effectors) changes its focus onto the $x$ and $y$ directions, producing lighter-colour heightmaps.
This may be caused by the fact that these losses distribute differently along the yield stress axis (the CD and EMD losses at level 1, the PRT CD loss at level 2). According to physics intuitions, a smaller yield stress will cause the object to yield more easily and respond more drastically to the poking forces, hence the more spreading in $x$ and $y$ directions and more compressing in the $z$ direction, and vice versa. One can see that a greater yield stress value in Table 3 corresponds to a darker heightmap in Figure 7, and vice versa.
Thirdly, though the particles and heightmaps are very similar, they are produced by quite different parameter combinations. For example, Table 3 reveals a positive correlation between the yield stress ($\boldsymbol{\sigma}_{y}$) and material density ($\rho$). One can see that, although the simulated results are similar, a larger $\boldsymbol{\sigma}_{y}$ is always accompanied by a larger $\rho$, and vice versa. This is realistic because when the object becomes heavier (greater $\rho$), it needs to be more difficult to yield; otherwise, it will collapse due to gravity. If the object is light (lower $\rho$), a smaller yield stress value allows it to exhibit the returning effect after compression.
In addition, from the level-2 results in Figure 7, one can observe another correlation between the friction coefficients and the material density. They are the three key parameters that determine how much the object will be moved in the shifting motion direction. The parameters in Table 3 and their corresponding visualisations show that,
@@
These relationships between parameters and visualisations align with our understanding of real-world physics, demonstrating that DPSI can produce physically plausible and interpretable parameter values.

### v2

@@ 3.4 Real-to-sim object reconstruction
Real-world platform: To collect real-world data, we set up a deformable object manipulation system where a Kuka IIWA LBR 14 industrial arm (Kuka, 2024) is equipped with a Zivid One+ medium camera (Zivid, 2024) for perceiving the object, and one of the three end-effectors, namely a rectangular cuboid, a cylinder roller and a bullet-shaped object for collecting interaction experiences with different contact geometries, as shown at the bottom-left of Figure 1. We use plasticine (non-hardening modelling clay) as the main material for our experiments.
@@
Reconstruction pipeline: To simulate the object as a set of particles, we design a pipeline to reconstruct the object particle system $\mathcal{X}^{rec}$ from the given fusion point cloud, as shown by the orange boxes in Figure 2. Firstly, we down-sample the processed fusion point cloud and create a watertight mesh using the ball-pivoting (BP) algorithm (Bernardini et al., 1999). The artificial bottom surface points prevent the BP algorithm from creating a non-flat bottom mesh surface. To sample particles, we create a set of voxels that fill in the reconstructed mesh surface uniformly given a resolution ($1080\times 1080$). Lastly, we fill particles into the voxels with a given density ($4\times 10^{6}$ particles$/m^{3}$). We assume that the particles are uniformly distributed within the mesh.
@@
Table 1: Waypoint designs and statistics of the interaction motions. For collecting optimisation data, there are two motions for each of the two levels of contact complexity. For out-of-distribution validation, one motion per end-effector is designed. Directions of the waypoints are relative to the robot base frame, as shown in Figure 1. Each waypoint of the real trajectories takes an uneven time interval, while each waypoint in simulation takes exactly $0.01$ second.
@@ 3.5 Real-to-sim trajectory reconstruction
All motions start from a configuration where the end-effector tip is positioned at the top centre of the object. ROS (Quigley et al., 2009) and the MoveIt!-based (Görner et al., 2019) OMPL planner (Sucan et al., 2012) are used to plan real-world motion trajectories $\tau^{real}$. For calculating each motion plan, we pass a series of waypoints to the MoveIt! planner by discretising each segment of the motion with an interval of $0.002$ m or $5^{\circ}$.
A challenging phenomenon that occurs during our data collection process is that the object tends to stick to the end-effectors after contact is made. If we allow this to happen and assume that the object always drops down eventually, the optimisation process will be extremely difficult due to a large uncertain dynamics process of the object dropping from the air. For example, the object may bounce out of the workspace or even off the table. This further exacerbates the difficulty of optimising with only the start and end observations of the object. However, it is beyond the capability of current perception hardware and thus a future research direction. To simplify the problem, we take a simple workaround by covering the end-effector with a thin layer of flour before each manipulation motion is executed. We sink the effector into the flour and flick it gently to ensure that it is only covered by a thin layer of flour. This greatly prevents the sticking phenomenon from happening. Note that the end-effector friction coefficient we are optimising for is then the one covered by flour instead of the original value.
Time duration constrained reconstruction: To simulate end-effector motions, the real-world trajectory generated by MoveIt! is inconvenient as it has uneven time differences between consecutive waypoints and the simulation can only handle a constant step size $dt$. Thus, for a motion segment between a pair of waypoints, we reconstruct its simulation counterpart to have constant velocity by dividing the travelled distance by the real-world time duration which is provided by the motion planner, and then we discretise the segment with a constant $dt$. In this study, we set $dt=0.01$ second for better simulation stability. A larger $dt$ will result in a too-high compounding error during simulation stepping, while a too-small $dt$ will demand too much computation. The statistics of the real and reconstructed motions are summarised in Table 1. It can be seen that the validation motions are much longer than the optimisation ones. The MoveIt! trajectories and the reconstructed ones are saved as .npy files and will be open to the public upon acceptance.
@@ 3.6 Loss functions
Finally, although the CD and EMD losses are very common in calculating point set distance, they are not intuitive to visualise and understand. Also, as revealed by the experiment results shown in Figures 4 and  5, the CD and EMD losses focus on different spatial aspects of two point sets, which make them biased for result analysis across all loss functions. To make results easier to analyse and compare in a more unbiased way, we further calculate the heightmaps of the fusion point cloud and the simulated particles, denoted as $\mathcal{I}_{T}^{real}$ and $\mathcal{I}_{T}^{sim}$. Specifically, we discretise a $0.11\times 0.11\text{m}^{2}$ square area centred at the centre of the object into a $32\times 32$ grid. For each cell of the grid, we find the point or particle in the x-y range with the highest $z$ value and assign that value as the height of that cell. To compare the distance between two heightmaps, we simply take their summed pixel difference as follows.
@@ 4.1 Experiment design
For each datapoint, the real-world object is initially roughly shaped into a convex shape. We then acquire its point cloud by fusing the captured multi-view data. Next, we calculate the coordinates of the top centre of the object, where the end-effector tip will be moved to, and the motion will always be executed from there. We perform the same point cloud capturing and fusion process after a motion is executed. Each datapoint takes no more than $2$ minutes to collect, with most of the time spent on capturing point clouds.
Optimisation & validation: At each contact level, for each pair of loss function and dataset, gradient descent is carried out with the Adam algorithm (Kingma and Ba, 2015) with three random seeds for $100$ iterations (gradient updates). For each datapoint, the simulation loads the initial particle configuration, simulates the motion and produces the resultant particle state, which is used to compute losses and gradients. Each optimisation iteration goes over all datapoints within a dataset and takes the average gradients to update the parameters. In-distribution validation is done after every gradient update with the validation dataset, simulating all datapoints and calculating the losses.
Table 2 displays the step sizes and the value ranges of the parameters of interest for gradient updates. To determine the appropriate step sizes, we randomly sample $20$ sets of parameter values and compute the statistical means and standard deviations of the gradients for each parameter based on the 12-mix dataset. We then randomly initialise the parameters, run the Adam optimiser for $100$ gradient updates with uniformly randomly generated gradients using the statistical means and deviations with different step sizes, and select the step size that allows the randomised parameters to converge in about $70$ gradient updates. The value ranges of the parameters are selected according to either the definitions of the physics models (Jones, 2009; Stomakhin et al., 2013) or values reported by various research on manual parameter identification experiments with similar real-world materials, such as soil and clay (StructX, 2014-2024; Waheed and Asmael, 2023). However, a too-high value of Young’s modulus $E$ was found to cause numerical instability in the simulation with the particles exploding away. Thus, an empirical upper bound for $E$ was determined by slowly increasing it until the simulation became unstable.
@@ 4.3 In-distribution performances
  • Does DPSI produce parameter values that are physically realistic and interpretable? Are they close to the parameter values reported in relevant literature and studies?
@@
However, the parameter distributions in Figure 5 clearly show that some parameters converge to much smaller and distinct value regions than those found in the level-1 experiments. More specifically, the values of Young’s modulus ($E$), Poisson’s ratio ($\nu$) and yield stress ($\boldsymbol{\sigma}_{y}$) become more converged, while they are more dispersed in the level-1 results. Interestingly, the material density ($\rho$) remains quite dispersed at both contact complexity levels. This is not surprising because, compared to the level-1 motions, the higher contact complexity induced by the shifting parts of the level-2 motions naturally poses a stricter selection range for the physics parameter values.
Also, some parameter values are closer to empirical studies of soft/hard/saturated clay in the level-2 experiments. For example, Young’s modulus values are closer to the reported range of $5000$ to $54000$ kPa, the material density seems to be closer to $1400$ kg/m3, and Poisson’s ratio is closer to the reported range of $0.4$ to $0.5$ (Waheed and Asmael, 2023; StructX, 2014-2024). However, the recovered yield stress values are higher than the reported values for clays ($210$ to $600$ kPa) (Rehman et al., 2018) because the reported values were found to make the material collapse in simulation. Based on the empirical values from Waheed and Asmael (2023); StructX (2014-2024) and the loss landscapes, we hand-pick a set of parameter values and calculate the validation losses at both levels. As denoted in Table 4, the resultant performances are reasonable but not as good as the performances of the DPSI method.
@@
$E$ (kPa) $\nu$ $\rho$ (kg/m3) $\boldsymbol{\sigma}_{y}$ (kPa) $\eta_{t}$ $\eta_{m}$
$30000$ $0.4$ $1330$ $1000$ $0.7$ $0.7$
Table 3: Hand-picked physics parameter values based on the loss landscapes, published literature and open-source website in geo-technical engineering (Rehman et al., 2018; Waheed and Asmael, 2023; StructX, 2014-2024). The yield stress value is slightly higher than the reported values because the reported values were found to cause the material collapse in simulation.
@@
*Validation height map (mm) with hand-picked parameters: 2122.59
@@
*Validation height map (mm) with hand-picked parameters: 2985.5688
@@
Table 4: The parameter values corresponding to the lowest validation heightmap distance in each optimisation case at both contact complexity levels. Bolded texts denote the three lowest losses at each level. It can be seen that the PRT EMD loss tends to produce the best results and DPSI can find parameters that outperform hand-picked parameters (from the geotechnical literature) with even just one interaction datapoint.
@@
Thirdly, the results show that it is possible to obtain comparable performances with just one datapoint at both contact complexity levels. This can be concluded by comparing the results horizontally: 1) most validation curves of the top five rows in Figs. 4 and 5 with 1-datapoint show highly similar tendencies with the 6-mix and 12-mix datasets, and 2) the last four rows in Figs. 4 and last six rows in Figure 5 also show that the parameters found with different datasets mostly converge to similar value regions. Moreover, Table 4 shows that the lowest validation heightmap distances at level-1 contact complexity and the third lowest at level-2 are achieved by optimising with the 1-rec. and 1-round datasets. These results are aligned with the analysis of the loss landscapes, where similar loss distributions are observed among different datasets.
(a) Level-1 contact, object configuration 1
(b) Level-1 contact, object configuration 2
(c) Level-2 contact, object configuration 1
(d) Level-2 contact, object configuration 2
@@
Figure 6: The particles and heightmaps of the objects after applying the second motions at contact level 1 (a-b) and 2 (c-d), simulated with the best set of physics parameters. Darker colours of the heightmaps indicate greater heights of the object. For each combination of the loss function and dataset at each level, three trajectories on two object configurations are simulated (3 effectors $\times$ 2 datapoints). There are three groups in each subfigure, with the ground truth point clouds and heightmaps placed in the leftmost column. In each group, a row shows the results corresponding to a loss function and a column results corresponding to a dataset.
@@
Qualitative results: We now examine the manipulation processes simulated using the best physics parameters corresponding to the lowest validation heightmap losses presented in Table 4. Figure 6 shows the resultant particles and heightmaps after the manipulation of objects at two contact complexity levels with three end-effectors (two object configurations are included for each contact level and end-effector).
@@
Secondly, by examining the details of the heightmaps more carefully, it shows that optimising the CD losses tends to produce heightmaps with darker colours (which represent greater heights of the object) while optimising the EMD losses tends to produce heightmaps where the objects look slightly bigger and wider. It indicates that optimising the CD losses tends to produce particles that match the height of the ground truth shapes, i.e., in the $z$ direction, while optimising the EMD losses tends to match the $x$ and $y$ directions. This is especially true at level-1 contact complexity, where the heightmaps in the first two rows for each object configuration in Subfigures 6(a) and 6(b) are generally darker than those in the last two rows. At level 2, the PRT CD loss sometimes changes its focus onto the $x$ and $y$ directions, producing lighter-colour heightmaps (the second rows of the heightmaps associated with the rectangle and round end-effectors in Subfigures 6(c) and 6(d)).
This may be caused by the fact that these losses distribute differently along the yield stress axis (the CD and EMD losses at level 1, the PRT CD loss at level 2). According to physics intuitions, a smaller yield stress will cause the object to yield more easily and respond more drastically to the poking forces, hence the more spreading in the $x$ and $y$ directions and more compressing in the $z$ direction, and vice versa. One can see that a greater yield stress value in Table 4 corresponds to a darker heightmap in Figure 6, and vice versa.
Thirdly, though the particles and heightmaps are very similar, they are produced by quite different parameter combinations. For example, Table 4 reveals a positive correlation between the yield stress ($\boldsymbol{\sigma}_{y}$) and material density ($\rho$). One can see that, although the simulated results are similar, a larger $\boldsymbol{\sigma}_{y}$ is always accompanied by a larger $\rho$, and vice versa. This is realistic because when the object becomes heavier (greater $\rho$), it needs to be more difficult to yield (deform plastically); otherwise, it will collapse due to gravity. If the object is light (lower $\rho$), a smaller yield stress value allows it to exhibit the returning effect after compression.
In addition, from the results at contact level 2 in Subfigures 6(c) and 6(d), one can also observe a correlation between the friction coefficients and the material density – the three key parameters that determine how much the object will be moved in the shifting motion direction. The parameters in Table 4 and their corresponding visualisations show that,
@@
These relationships between parameters and visualisations align with our understanding of real-world physics, demonstrating that DPSI can produce physically plausible and interpretable parameter values. Moreover, the variations in parameter values suggest that, aside from the mismatch of the models and the materials, there is hardly any fixed set of parameter values that can simulate a material in every condition. These parameters will always have some differences in different conditions (e.g., heat, light, humidity, etc.) and they change over time (e.g., drying). Therefore, the fact that DPSI can estimate the parameters with only one interaction makes it highly promising for fast parameter re-identification that can be performed online to recalibrate simulation accuracy efficiently, eliminating the need to keep track of previously identified parameter values.